OpenStudy (anonymous):

Evaluate x/y,for x=2/3 and y= 5/6 A. 5/9 B. 4/5 C 5/4 D 9/5

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

I thinks its C

OpenStudy (skullpatrol):

Why?

OpenStudy (hba):

Show me your working @JhonDoe

OpenStudy (skullpatrol):



OpenStudy (anonymous):

NVM its 9/5 because find the common denaminater wich is 12 and you add them then you simplify them witch its 9/5

OpenStudy (skullpatrol):



OpenStudy (hba):

You are still wrong and show us your working @JhonDoe

OpenStudy (anonymous):

ohh @skullpatrol but how do two fractions make a fraction?

OpenStudy (skullpatrol):

You have to simplify it.

OpenStudy (anonymous):

4/5, \[\frac{ \frac{ 2 }{ 3 } }{ \frac{ 5 }{ 6 } } = \frac{ 2*6 }{ 3*5 } = \frac{ 12 }{ 15 = \frac{ 4 }{ 5 } }\]

OpenStudy (anonymous):

the final equals on the bottom of that fraction was supposed to be 12/15=4/5

OpenStudy (hba):

According to a rule \[\huge\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b } \times \frac{ d }{ c }\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[\left(\begin{matrix}2/3 \\ 5/6\end{matrix}\right)=\frac{ 2*6 }{ 3*5}=\frac{ 12}{15} = \frac{ 4}{5 }\]

OpenStudy (anonymous):

@skullpatrol